# Graph Laplacian-based Bayesian Multi-fidelity Modeling

### v1

@@ 2.3 Selection of high-fidelity data
In this section, we describe the policy used to select $M\ll N$ high-fidelity data points to acquire. This is accomplished by utilizing the graph constructed from the low-fidelity data to determine $M$ clusters and their corresponding centroids. Thereafter, the data points closest to each of these centroids are determined, and the corresponding high-fidelity data is acquired. The logic in selecting these points is driven by the observation that points that belong to a given cluster will tend to move together from their low-fidelity coordinates to their high-fidelity coordinates. Thus the displacement vector for the centroid is likely to be representative of the displacement vector for most of the points in the cluster. The steps in this process are:
@@
    Use $k$-means clustering on the points $\{\bm{\xi}^{(i)}\}_{i=1}^{N}$ to find $M$ clusters.
@@ 2.4 Construction of the likelihood
$\hat{\bm{\Phi}}:=\begin{pmatrix}\hat{\bm{\phi}}^{(1)}\\ \vdots\\ \hat{\bm{\phi}}^{(M)}\end{pmatrix}=\begin{pmatrix}\hat{\bm{\phi}}_{1}&\cdots&\hat{\bm{\phi}}_{D}\end{pmatrix}\in\mathbb{R}^{M\times D}.$
@@
$\hat{\bm{u}}^{(i)}=\bm{u}^{(i)}+\bm{\varepsilon}^{(i)},\qquad i\in\{1,\,\dots,\,M\},$
@@
$\hat{\bm{\phi}}^{(i)}=\bm{\phi}^{(i)}+\bm{\varepsilon}^{(i)},\qquad i\in\{1,\,\dots,\,M\}.$
@@
$p(\hat{\bm{\Phi}}|\bm{\Phi})\propto\exp\left(-\frac{\|\hat{\bm{\Phi}}-\bm{P}_{M}\bm{\Phi}\|_{F}^{2}}{2\sigma^{2}}\right),$ (2.6)
@@ 2.5 Definition of the posterior density
$\begin{split}p(\bm{\Phi}|\hat{\bm{\Phi}})&\propto p(\hat{\bm{\Phi}}|\bm{\Phi})p(\bm{\Phi})\\ &\propto\exp\left(-\frac{\|\hat{\bm{\Phi}}-\bm{P}_{M}\bm{\Phi}\|_{F}^{2}}{2\sigma^{2}}-\frac{\omega}{2}\langle\bm{\Phi},\left(\bm{L}+\tau\bm{I}_{N}\right)^{\beta}\bm{\Phi}\rangle_{F}\right)\\ &=\prod_{m=1}^{D}\exp\left(-\frac{\|\hat{\bm{\phi}}_{m}-\bm{P}_{M}\bm{\phi}_{m}\|_{2}^{2}}{2\sigma^{2}}-\frac{\omega}{2}\bm{\phi}_{m}^{T}\left(\bm{L}+\tau\bm{I}_{N}\right)^{\beta}\bm{\phi}_{m}\right),\end{split}$ (2.7)
where we have made use of (2.4) and (2.6). Since both the prior and the likelihood are normal distributions, we can write the expression above as
$\begin{split}p(\bm{\Phi}|\hat{\bm{\Phi}})&\propto\exp\left(-\frac{1}{2}\langle\bm{\Phi}-\bm{\Phi}^{*},\bm{C}^{-1}(\bm{\Phi}-\bm{\Phi}^{*})\rangle_{F}\right)\\ &=\prod_{m=1}^{D}\exp\left(-\frac{1}{2}\left(\bm{\phi}_{m}-\bm{\phi}^{*}_{m}\right)^{T}\cdot\bm{C}^{-1}\left(\bm{\phi}_{m}-\bm{\phi}^{*}_{m}\right)\right),\end{split}$ (2.8)
@@
$\bm{C}$ $=\left(\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\bm{P}_{M}+\omega(\bm{L}+\tau\bm{I}_{N})^{\beta}\right)^{-1},$ (2.9)
$\bm{\Phi}^{*}$ $=\frac{1}{\sigma^{2}}\bm{C}\bm{P}_{M}^{T}\hat{\bm{\Phi}}.$ (2.10)
@@
It is instructive to write (2.10) as a solution to a linear system of equations. This yields,
$\left(\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\bm{P}_{M}+\omega(\bm{L}+\tau\bm{I}_{N})^{\beta}\right)\bm{\Phi}^{*}=\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\hat{\bm{\Phi}}.$ (2.11)
@@ 2.6 Selection of the hyper-parameters
The method described requires the evaluation of the three hyper-parameters $\tau$, $\omega$ and $\beta$. The parameter $\tau>0$ was artificially introduced into the problem to guarantee invertibility of $\bm{L}+\tau\bm{I}_{N}$, and therefore allows the interpretation of our set-up as a Bayesian inverse problem with (2.9) providing the expression of the covariance matrix of the posterior distribution. As described in Section 2.2.1, $\tau$ is set equal to the smallest non-zero eigenvalue of the graph Laplacian.
Next, we comment on the role of the regularization exponent $\beta$. When maximizing the posterior distribution $p(\bm{\Phi}|\hat{\bm{\Phi}})$ over $\bm{\Phi}$, higher values of $\beta$ enforce that the resulting $\bm{\phi}_{m}$ has smooth variation within clusters also for higher derivatives; in short, it enforces control of derivatives of order $\beta$ and may be selected $\beta\geq 1$ [40]. Given this, we select $\beta=2$ and note that this choice is adequate in all our numerical experiments.
@@
$\frac{1}{N-M}\sum_{i=M}^{N}\sqrt{C_{ii}}=r\sigma,$
for some $r>1$. This guarantees that the confidence of the multi-fidelity model is not greater than the one of the high-fidelity model. In our numerical experiments we have observed that that a value of $r=3$ leads to good results for all cases. Remarkably, the selection of the hyper-parameters using the approach described above does not require knowledge of any additional high-fidelity data. All the high-fidelity data is used to estimate the mode of the true data points.
@@ 3.1 Expansion in a truncated eigenfunction basis
$a:=2\max_{i\in\{1,...,N\}}D_{ii}^{1-p-q}.$
Then it is a standard result that the eigenvalues of $\bm{L}$ lie in $[0,a]$ [49, Lemma 2.5(f)]. [^2] To apply this Lemma, we use that $\bm{L}$ is similar to $\bm{D}^{-(p+q)}(\bm{D}-\bm{W})$. Next, compute the $K$ leading eigenvalues $\lambda^{\prime}_{k}$ and eigenvectors ${\bm{\psi}}^{\prime}_{k}$ of the symmetric positive semi-definite matrix $a\bm{I}_{N}-\bm{L}$. Finally, the low-lying eigenvectors of $\bm{L}$ are $\bm{\psi}_{k}=\bm{\psi}^{\prime}_{k}$, with corresponding eigenvalues $\lambda_{k}=a-\lambda^{\prime}_{k}$.
@@
$\bm{\Psi}_{K}$ $=\begin{pmatrix}\bm{\psi}_{1}&\cdots&\bm{\psi}_{K}\end{pmatrix}\in\mathbb{R}^{N\times K},$
$\bm{\Lambda}_{K}$ $=\operatorname{diag}\left(\lambda_{1},\,\dots,\,\lambda_{K}\right)\in\mathbb{R}^{K\times K},$
@@
$\bm{\phi}_{m}=\sum_{k=1}^{K}A_{km}\bm{\psi}_{k}.$
Using this expansion in the posterior distribution for the displacement (2.7), we arrive at an expression for the distribution for the coefficients $\bm{A}_{K}$,
$p(\bm{A}_{K}|\hat{\bm{\Phi}})$ $\propto\exp\left(-\frac{\|\hat{\bm{\Phi}}-\bm{P}_{M}\bm{\Psi}_{K}\bm{A}_{K}\|_{F}^{2}}{2\sigma^{2}}-\frac{\omega}{2}\operatorname{tr}\left(\bm{A}_{K}^{T}\bm{\Psi}_{K}^{T}\left(\bm{L}+\tau\bm{I}_{N}\right)^{\beta}\bm{\Psi}_{K}\bm{A}_{K}\right)\right)$
$=\exp\left(-\frac{\|\hat{\bm{\Phi}}-\bm{P}_{M}\bm{\Psi}_{K}\bm{A}_{K}\|_{F}^{2}}{2\sigma^{2}}-\frac{\omega}{2}\operatorname{tr}\left(\bm{A}_{K}^{T}\left(\bm{\Lambda}_{K}+\tau\bm{I}_{K}\right)^{\beta}\bm{A}_{K}\right)\right),$ (3.1)
@@
Recognizing that the distribution for $\bm{A}_{K}$ is the product of two normal distributions and is therefore itself a normal distribution, we may write the distribution Equation 3.1 as,
$p(\bm{A}_{K}|\hat{\bm{\Phi}})$ $\propto\exp\left(\frac{1}{2}\langle\bm{A}_{K}-\bm{A}_{K}^{*},\bm{C}_{A_{K}}^{-1}(\bm{A}_{K}-\bm{A}_{K}^{*})\rangle_{F}\right)$
$=\prod_{m=1}^{D}\exp\left(-\frac{1}{2}\left(\bm{a}_{m}-\bm{a}^{*}_{m}\right)^{T}\cdot\bm{C}_{A_{K}}^{-1}\left(\bm{a}_{m}-\bm{a}^{*}_{m}\right)\right)$
@@
$\bm{C}_{A_{K}}^{-1}$ $=\frac{1}{\sigma^{2}}\left(\bm{P}_{M}\bm{\Psi}_{K}\right)^{T}\bm{P}_{M}\bm{\Psi}_{K}+\omega(\bm{\Lambda}_{K}+\tau\bm{I}_{K})^{\beta},$ (3.2)
$\bm{A}_{K}^{*}$ $=\frac{1}{\sigma^{2}}\bm{C}_{A_{K}}\left(\bm{P}_{M}\bm{\Psi}_{K}\right)^{T}\hat{\bm{\Phi}}.$ (3.3)
@@ 3.1.1 Computational complexities
The quantities to be computed and stored for this method, and the computational complexity of doing so, are summarized in Table 1. For the complexity of computing the $K$ leading eigenvalues and eigenvectors of $a\bm{I}_{N}-\bm{L}$ (which is the bottleneck for computing the truncated eigendecomposition), we assume that randomized block Krylov methods [50] are used.
@@
$\bm{\Psi}_{K},\bm{\Lambda}_{K}$ $\mathcal{O}(N^{2}K\log N)$ [50] $\mathcal{O}(N^{2}(\log N)^{2})$ [50]
@@ 3.2 Low rank Nyström approximation
Recall that to compute $\bm{\Phi}^{*}$ we need to solve (2.11). In this section we describe a way to compute this in a way that is very efficient in $N$, in the setting where $\mathcal{O}(N^{2})$ time and space complexity are both intractable. We can do this in the special case of $p=q=1/2$ by computing a low rank approximation of $(\bm{L}+\tau\bm{I}_{N})^{\beta}$ via the Nyström extension.
@@ 3.2.1 Nyström-QR approximation
[…]
$\bm{W}\approx\bm{W}(:,X)\bm{W}(X,X)^{\dagger}\bm{W}(X,:),$ (3.4)
where we have used the Matlab notation $\bm{W}(:,X):=[W_{ij}]_{i=1,j\in X}^{N}$ and $\bm{W}(X,X):=[W_{ij}]_{i,j\in X}$, and $\bm{W}(X,X)^{\dagger}$ denotes the pseudoinverse of $\bm{W}(X,X)$. Observe that this approximation of $\bm{W}$ has at most rank $K$. In this subsection we will use this approximation to derive a low-rank approximation for $(\bm{L}+\tau\bm{I}_{N})^{\beta}$, and thereby efficiently approximate solutions to Equation 2.11 and matrix-vector products with $\bm{C}$.
@@
The Nyström extension is most well-behaved applied to symmetric positive semidefinite matrices, as if $\bm{S}$ is such a matrix then the condition number of $\bm{S}(X,X)$ is bounded above by that of $\bm{S}$. However, this is not satisfied by $\bm{W}$, which is constructed to have zeroes on the diagonal (i.e., the graph has no self-loops) and hence has trace zero, and is therefore an indefinite matrix. In the indefinite case, positive and negative eigenvalues can combine to leave $\bm{W}(X,X)$ with much smaller (in magnitude) non-zero eigenvalues than $\bm{W}$ and correspondingly a much higher condition number. This case has seen recent attention in Nakatsukasa and Park [54], which recommends taking $K$ higher than one’s desired rank $r$, and then truncating the spectrum of $\bm{W}(X,X)$ to the $r$ largest (in magnitude) eigenvalues before computing the pseudoinverse, to control the condition number. We will ignore this refinement in the below, but it is only a minor modification to include.
Given Equation 3.4, we can therefore compute
$\bm{D}\approx\hat{\bm{D}}:=\operatorname{diag}(\bm{W}(:,X)\bm{W}(X,X)^{\dagger}\bm{W}(X,:)\mathbf{1}_{N})$
@@
$\bm{L}\approx\bm{I}_{N}-\hat{\bm{D}}^{-\frac{1}{2}}\bm{W}(:,X)\bm{W}(X,X)^{\dagger}\bm{W}(X,:)\hat{\bm{D}}^{-\frac{1}{2}}.$
Next, we can extract an approximate low rank eigendecomposition of $\bm{L}$ from this approximation by following a method of Bebendorf and Kunis [55], first recommended in the graph-based learning context by Alfke et al. [56]. We compute the thin QR factorization
$\bm{Q}\bm{R}=\hat{\bm{D}}^{-\frac{1}{2}}\bm{W}(:,X),$
@@
$\bm{R}\bm{W}(X,X)^{\dagger}\bm{R}^{T}=\bm{\Gamma}\bm{\Sigma}\bm{\Gamma}^{T},$
@@
$\bm{L}\approx\bm{I}_{N}-\tilde{\bm{U}}\bm{\Sigma}\tilde{\bm{U}}^{T},$
@@
$(\bm{L}+\tau\bm{I}_{N})^{\beta}$ $\approx((1+\tau)\bm{I}_{N}-\tilde{\bm{U}}\bm{\Sigma}\tilde{\bm{U}}^{T})^{\beta}$
$=\left((1+\tau)(\bm{I}_{N}-\tilde{\bm{U}}\tilde{\bm{U}}^{T})+\tilde{\bm{U}}\left((1+\tau)\bm{I}_{K}-\bm{\Sigma}\right)\tilde{\bm{U}}^{T}\right)^{\beta}$
$=(1+\tau)^{\beta}(\bm{I}_{N}-\tilde{\bm{U}}\tilde{\bm{U}}^{T})^{\beta}+\left(\tilde{\bm{U}}((1+\tau)\bm{I}_{K}-\bm{\Sigma})\tilde{\bm{U}}^{T}\right)^{\beta}$ (*)
$=(1+\tau)^{\beta}(\bm{I}_{N}-\tilde{\bm{U}}\tilde{\bm{U}}^{T})+\tilde{\bm{U}}((1+\tau)\bm{I}_{K}-\bm{\Sigma})^{\beta}\tilde{\bm{U}}^{T}$ (**∗ ∗)
$=(1+\tau)^{\beta}\bm{I}_{N}+\tilde{\bm{U}}\left(((1+\tau)\bm{I}_{K}-\bm{\Sigma})^{\beta}-(1+\tau)^{\beta}\bm{I}_{K}\right)\tilde{\bm{U}}^{T}.$
Here line ()(*)( ∗ ) follows from the fact that if $\bm{AB}=\bm{BA}=\bm{0}$ and $\bm{A},\bm{B}$ are diagonalizable, then $(\bm{A}+\bm{B})^{\beta}=\bm{A}^{\beta}+\bm{B}^{\beta}$, which is here satisfied by $\bm{A}=(1+\tau)(\bm{I}_{N}-\tilde{\bm{U}}\tilde{\bm{U}}^{T})$ and $\bm{B}=\tilde{\bm{U}}((1+\tau)\bm{I}_{K}-\bm{\Sigma})\tilde{\bm{U}}^{T}$, since $\tilde{\bm{U}}^{T}\tilde{\bm{U}}=\bm{I}_{K}$ and $\bm{A}$ and $\bm{B}$ are symmetric matrices. Line ()(**)( ∗ ∗ ) follows from the fact that $\tilde{\bm{U}}^{T}\tilde{\bm{U}}=\bm{I}_{K}$ and therefore that $\bm{I}_{N}-\tilde{\bm{U}}\tilde{\bm{U}}^{T}$ is idempotent.
3.2.2 Solving (2.11)

### v2

@@ 2.3 Selection of high-fidelity data
In this section, we describe the policy used to select $M\ll N$ high-fidelity data points to acquire. This is accomplished by performing spectral clustering on the graph constructed from the low-fidelity data to determine $M$ clusters and their corresponding centroids. Thereafter, the data points closest to each of these centroids are determined, and the corresponding high-fidelity data is acquired. The logic of using spectral clustering is that it is consistent with our choice of prior, based on a graph Laplacian regularizer. Whereas, the choice of selecting the centroids is driven by the observation that points that belong to a given cluster will tend to move together from their low-fidelity coordinates to their high-fidelity coordinates. Thus the displacement vector for the centroid is likely to be representative of the displacement vector for most of the points in the cluster. The steps in this process are:
@@
    Perform clustering on the points $\{\bm{\xi}^{(i)}\}_{i=1}^{N}$ to find $M$ clusters. For this task, standard clustering techniques such as $k$-means or DBSCAN can be used. [^4] In this work, we assume that the number $M$ of computed eigenfunctions equals the number of desired high-fidelity points, and therefore of clusters. However, if one desired to decouple these quantities, then an advantage of DBSCAN is that it does not require a pre-specified number of clusters, unlike $k$-means.
@@ 2.4 Construction of the likelihood
$\hat{\bm{\Phi}}:=\begin{pmatrix}\hat{\bm{\phi}}^{(1)}\\ \vdots\\ \hat{\bm{\phi}}^{(M)}\end{pmatrix}=\begin{pmatrix}\hat{\bm{\phi}}_{1}&\cdots&\hat{\bm{\phi}}_{D}\end{pmatrix}\in\mathbb{R}^{M\times D}.$ (11)
@@
$\hat{\bm{u}}^{(i)}=\bm{u}^{(i)}+\bm{\varepsilon}^{(i)},\qquad i\in\{1,\,\dots,\,M\},$ (12)
@@
$\hat{\bm{\phi}}^{(i)}=\bm{\phi}^{(i)}+\bm{\varepsilon}^{(i)},\qquad i\in\{1,\,\dots,\,M\}.$ (13)
@@
$p(\hat{\bm{\Phi}}|\bm{\Phi})\propto\exp\left(-\frac{\|\hat{\bm{\Phi}}-\bm{P}_{M}\bm{\Phi}\|_{F}^{2}}{2\sigma^{2}}\right),$ (14)
@@ 2.5 Definition of the posterior density
$\begin{split}p(\bm{\Phi}|\hat{\bm{\Phi}})&\propto p(\hat{\bm{\Phi}}|\bm{\Phi})p(\bm{\Phi})\\ &\propto\exp\left(-\frac{\|\hat{\bm{\Phi}}-\bm{P}_{M}\bm{\Phi}\|_{F}^{2}}{2\sigma^{2}}-\frac{\omega}{2}\langle\bm{\Phi},\left(\bm{L}+\tau\bm{I}_{N}\right)^{\beta}\bm{\Phi}\rangle_{F}\right)\\ &=\prod_{m=1}^{D}\exp\left(-\frac{\|\hat{\bm{\phi}}_{m}-\bm{P}_{M}\bm{\phi}_{m}\|_{2}^{2}}{2\sigma^{2}}-\frac{\omega}{2}\bm{\phi}_{m}^{T}\left(\bm{L}+\tau\bm{I}_{N}\right)^{\beta}\bm{\phi}_{m}\right),\end{split}$ (16)
where we have made use of (8) and (14). Since both the prior and the likelihood are normal distributions, we can write the expression above as
$\begin{split}p(\bm{\Phi}|\hat{\bm{\Phi}})&\propto\exp\left(-\frac{1}{2}\langle\bm{\Phi}-\bm{\Phi}^{*},\bm{C}^{-1}(\bm{\Phi}-\bm{\Phi}^{*})\rangle_{F}\right)\\ &=\prod_{m=1}^{D}\exp\left(-\frac{1}{2}\left(\bm{\phi}_{m}-\bm{\phi}^{*}_{m}\right)^{T}\cdot\bm{C}^{-1}\left(\bm{\phi}_{m}-\bm{\phi}^{*}_{m}\right)\right),\end{split}$ (17)
@@
$\bm{C}$ $=\left(\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\bm{P}_{M}+\omega(\bm{L}+\tau\bm{I}_{N})^{\beta}\right)^{-1},$ (18)
$\bm{\Phi}^{*}$ $=\frac{1}{\sigma^{2}}\bm{C}\bm{P}_{M}^{T}\hat{\bm{\Phi}}$. (19)
@@
It is instructive to write (19) as a solution to a linear system of equations. This yields,
$\left(\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\bm{P}_{M}+\omega(\bm{L}+\tau\bm{I}_{N})^{\beta}\right)\bm{\Phi}^{*}=\frac{1}{\sigma^{2}}\bm{P}_{M}^{T}\hat{\bm{\Phi}}$. (20)
@@ 2.6 Selection of the hyper-parameters
The method described requires the evaluation of the three hyper-parameters $\tau$, $\omega$ and $\beta$. The parameter $\tau>0$ was artificially introduced into the problem to guarantee invertibility of $\bm{L}+\tau\bm{I}_{N}$, and therefore allows the interpretation of our set-up as a Bayesian inverse problem with (18) providing the expression of the covariance matrix of the posterior distribution. As described in Section 2.2.1, $\tau$ is set equal to the smallest non-zero eigenvalue of the graph Laplacian.
[…]
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{N}\sum_{i=1}^{N}\sqrt{C_{ii}}=r\sigma,$} (21)
for some $r>1$. This guarantees that the confidence of the multi-fidelity model is not greater than the one of the high-fidelity model. In our numerical experiments we have observed that that a value of $r=3$ leads to good results for all cases. Remarkably, the selection of the hyper-parameters using the approach described above does not require knowledge of any additional high-fidelity data. All the high-fidelity data is used to estimate the mode of the true data points. Additionally, in B we show that the gradient of the function
$\mathcal{J}(\omega):=\left(\frac{1}{N}\sum_{i=1}^{N}\sqrt{C_{ii}(\omega)}-r\sigma\right)^{2},$ (22)
which needs to be minimized to solve (21), can be found in closed form.
@@ 3.1 Expansion in a truncated eigenfunction basis
$a:=2\max_{i\in\{1,...,N\}}D_{ii}^{1-p-q}.$ (23)
Then it is a standard result that the eigenvalues of $\bm{L}$ lie in $[0,a]$ [vanGennip14, Lemma 2.5(f)]. [^5] To apply this Lemma, we use that $\bm{L}$ is similar to $\bm{D}^{-(p+q)}(\bm{D}-\bm{W})$. Next, compute the $K$ leading eigenvalues $\lambda^{\prime}_{k}$ and eigenvectors ${\bm{\psi}}^{\prime}_{k}$ of the symmetric positive semi-definite matrix $a\bm{I}_{N}-\bm{L}$. Finally, the low-lying eigenvectors of $\bm{L}$ are $\bm{\psi}_{k}=\bm{\psi}^{\prime}_{k}$, with corresponding eigenvalues $\lambda_{k}=a-\lambda^{\prime}_{k}$.
Remark.
Choosing smaller values of the cutoff value $K$ reduces the computational budget of the method (see Table 1). In practice, $K$ is a parameter that is often fixed using heuristics. In preliminary experiments (a more thorough investigation is the subject of ongoing work) we observed non-monotonic error dependence on $K$, often seeing a double-descent phenomenon. It is thus hard to recommend principled heuristics for $K$, except to note that when we did observe a second descent in error as a function of $K$, it roughly was for $K>M$, the number of high-fidelity points. Alternatively, $K$ can be tuned as a parameter using a holdout validation set, wherein a $K$ is sought (within computational budget) which minimizes validation error.
@@
$\bm{\Psi}_{K}$ $=\begin{pmatrix}\bm{\psi}_{1}&\cdots&\bm{\psi}_{K}\end{pmatrix}\in\mathbb{R}^{N\times K},$ (24)
$\bm{\Lambda}_{K}$ $=\operatorname{diag}\left(\lambda_{1},\,\dots,\,\lambda_{K}\right)\in\mathbb{R}^{K\times K},$ (25)
@@
$\bm{\phi}_{m}=\sum_{k=1}^{K}A_{km}\bm{\psi}_{k}.$ (27)
Using this expansion in the posterior distribution for the displacement (16), we arrive at an expression for the distribution for the coefficients $\bm{A}_{K}$,
[…]
$=\exp\left(-\frac{\|\hat{\bm{\Phi}}-\bm{P}_{M}\bm{\Psi}_{K}\bm{A}_{K}\|_{F}^{2}}{2\sigma^{2}}-\frac{\omega}{2}\operatorname{tr}\left(\bm{A}_{K}^{T}\left(\bm{\Lambda}_{K}+\tau\bm{I}_{K}\right)^{\beta}\bm{A}_{K}\right)\right),$ (28)
@@
Recognizing that the distribution for $\bm{A}_{K}$ is the product of two normal distributions and is therefore itself a normal distribution, we may write the distribution Equation 28 as,
[…]
$=\prod_{m=1}^{D}\exp\left(-\frac{1}{2}\left(\bm{a}_{m}-\bm{a}^{*}_{m}\right)^{T}\cdot\bm{C}_{A_{K}}^{-1}\left(\bm{a}_{m}-\bm{a}^{*}_{m}\right)\right),$ (29)
@@
$\bm{C}_{A_{K}}^{-1}$ $=\frac{1}{\sigma^{2}}\left(\bm{P}_{M}\bm{\Psi}_{K}\right)^{T}\bm{P}_{M}\bm{\Psi}_{K}+\omega(\bm{\Lambda}_{K}+\tau\bm{I}_{K})^{\beta},$ (30)
$\bm{A}_{K}^{*}$ $=\frac{1}{\sigma^{2}}\bm{C}_{A_{K}}\left(\bm{P}_{M}\bm{\Psi}_{K}\right)^{T}\hat{\bm{\Phi}}$. (31)
@@ 3.1.1 Computational complexities
The quantities to be computed and stored for this method, and the computational complexity of doing so, are summarized in Table 1. For the complexity of computing the $K$ leading eigenvalues and eigenvectors of $a\bm{I}_{N}-\bm{L}$ (which is the bottleneck for computing the truncated eigendecomposition), we assume that randomized block Krylov methods [MuscoMusco15] are used.
@@
$\bm{\Psi}_{K},\bm{\Lambda}_{K}$ $\mathcal{O}(N^{2}K\log N)$ [MuscoMusco15] $\mathcal{O}(N^{2}(\log N)^{2})$ [MuscoMusco15]
@@ 3.2 Low rank Nyström approximation
Recall that to compute $\bm{\Phi}^{*}$ we need to solve (20). In this section we describe a way to compute this in a way that is very efficient in $N$, in the setting where $\mathcal{O}(N^{2})$ time and space complexity are both intractable. We can do this in the special case of $p=q=1/2$ by computing a low rank approximation of $(\bm{L}+\tau\bm{I}_{N})^{\beta}$ via the Nyström extension.
@@ 3.2.1 Nyström-QR approximation
Notice that, since $p=q=1/2$, $\bm{L}:=\bm{I}_{N}-\bm{D}^{-1/2}\bm{W}\bm{D}^{-1/2}$. Choosing $X\subseteq\{1,...,N\}$ at random (with perhaps the condition that $X$ contain some or all of the high-fidelity points) with $|X|=:K$, we can approximate $\bm{W}$ via the Nyström extension [Nystrom30, FowlkesBelongieChungMalik04, BelongieFowlkesChungMalik02]:
$\bm{W}\approx\bm{W}(:,X)\bm{W}(X,X)^{\dagger}\bm{W}(X,:),$ (32)
where we have used the Matlab notation $\bm{W}(:,X):=[W_{ij}]_{i=1,j\in X}^{N}$ and $\bm{W}(X,X):=[W_{ij}]_{i,j\in X}$, and $\bm{W}(X,X)^{\dagger}$ denotes the pseudoinverse of $\bm{W}(X,X)$. Observe that this approximation of $\bm{W}$ has at most rank $K$. In this subsection we will use this approximation to derive a low-rank approximation for $(\bm{L}+\tau\bm{I}_{N})^{\beta}$, and thereby efficiently approximate solutions to Equation 20 and matrix-vector products with $\bm{C}$.
@@
The Nyström extension is most well-behaved applied to symmetric positive semidefinite matrices, as if $\bm{S}$ is such a matrix then the condition number of $\bm{S}(X,X)$ is bounded above by that of $\bm{S}$. However, this is not satisfied by $\bm{W}$, which is constructed to have zeroes on the diagonal (i.e., the graph has no self-loops) and hence has trace zero, and is therefore an indefinite matrix. In the indefinite case, positive and negative eigenvalues can combine to leave $\bm{W}(X,X)$ with much smaller (in magnitude) non-zero eigenvalues than $\bm{W}$ and correspondingly a much higher condition number. This case has seen recent attention in Nakatsukasa and Park [NakatsukasaPark23], which recommends taking $K$ higher than one’s desired rank $r$, and then truncating the spectrum of $\bm{W}(X,X)$ to the $r$ largest (in magnitude) eigenvalues before computing the pseudoinverse, to control the condition number. We will ignore this refinement in the below, but it is only a minor modification to include.
Given Equation 32, we can therefore compute
$\bm{D}\approx\hat{\bm{D}}:=\operatorname{diag}(\bm{W}(:,X)\bm{W}(X,X)^{\dagger}\bm{W}(X,:)\mathbf{1}_{N})$ (33)
@@
$\bm{L}\approx\bm{I}_{N}-\hat{\bm{D}}^{-\frac{1}{2}}\bm{W}(:,X)\bm{W}(X,X)^{\dagger}\bm{W}(X,:)\hat{\bm{D}}^{-\frac{1}{2}}.$ (34)
Next, we can extract an approximate low rank eigendecomposition of $\bm{L}$ from this approximation by following a method of Bebendorf and Kunis [BebendorfKunis09], first recommended in the graph-based learning context by Alfke et al. [AlfkePottsStollVolkmer18]. We compute the thin QR factorization
$\bm{Q}\bm{R}=\hat{\bm{D}}^{-\frac{1}{2}}\bm{W}(:,X),$ (35)
@@
$\bm{R}\bm{W}(X,X)^{\dagger}\bm{R}^{T}=\bm{\Gamma}\bm{\Sigma}\bm{\Gamma}^{T},$ (36)
@@
$\bm{L}\approx\bm{I}_{N}-\tilde{\bm{U}}\bm{\Sigma}\tilde{\bm{U}}^{T},$ (37)
@@
$(\bm{L}+\tau\bm{I}_{N})^{\beta}$ $\approx((1+\tau)\bm{I}_{N}-\tilde{\bm{U}}\bm{\Sigma}\tilde{\bm{U}}^{T})^{\beta}$ (38)
$=\left((1+\tau)(\bm{I}_{N}-\tilde{\bm{U}}\tilde{\bm{U}}^{T})+\tilde{\bm{U}}\left((1+\tau)\bm{I}_{K}-\bm{\Sigma}\right)\tilde{\bm{U}}^{T}\right)^{\beta}$ (39)
$=(1+\tau)^{\beta}(\bm{I}_{N}-\tilde{\bm{U}}\tilde{\bm{U}}^{T})^{\beta}+\left(\tilde{\bm{U}}((1+\tau)\bm{I}_{K}-\bm{\Sigma})\tilde{\bm{U}}^{T}\right)^{\beta}$ (40)
$=(1+\tau)^{\beta}(\bm{I}_{N}-\tilde{\bm{U}}\tilde{\bm{U}}^{T})+\tilde{\bm{U}}((1+\tau)\bm{I}_{K}-\bm{\Sigma})^{\beta}\tilde{\bm{U}}^{T}$ (41)
$=(1+\tau)^{\beta}\bm{I}_{N}+\tilde{\bm{U}}\left(((1+\tau)\bm{I}_{K}-\bm{\Sigma})^{\beta}-(1+\tau)^{\beta}\bm{I}_{K}\right)\tilde{\bm{U}}^{T}.$ (42)
Here line (40) follows from the fact that if $\bm{AB}=\bm{BA}=\bm{0}$ and $\bm{A},\bm{B}$ are diagonalizable, then $(\bm{A}+\bm{B})^{\beta}=\bm{A}^{\beta}+\bm{B}^{\beta}$, which is here satisfied by $\bm{A}=(1+\tau)(\bm{I}_{N}-\tilde{\bm{U}}\tilde{\bm{U}}^{T})$ and $\bm{B}=\tilde{\bm{U}}((1+\tau)\bm{I}_{K}-\bm{\Sigma})\tilde{\bm{U}}^{T}$, since $\tilde{\bm{U}}^{T}\tilde{\bm{U}}=\bm{I}_{K}$ and $\bm{A}$ and $\bm{B}$ are symmetric matrices. Line (41) follows from the fact that $\tilde{\bm{U}}^{T}\tilde{\bm{U}}=\bm{I}_{K}$ and therefore that $\bm{I}_{N}-\tilde{\bm{U}}\tilde{\bm{U}}^{T}$ is idempotent.
3.2.2 Solving (20)